OpenStudy (anonymous):

Find the average value of the function f(x) = (x + 3)^2 on the interval [0, 1]

OpenStudy (anonymous):

integrate first to find area, then multiply that number by (1)/(1-0)

OpenStudy (anonymous):

The integral is: \[\frac{1}{b-a}\int\limits_{\alpha}^{\beta}f(x)dx\] Where the interval is: \[[a,b]\] In this case you have [0,1] so that makes the interval length 1. So, in this case, the average value is the area. So your integral becomes: \[\int\limits_0^1 (x+3)^2 dx\] \[\frac{(x+3)^3}{3}|_0^1=\frac{4^3}{3}-\frac{3^3}{3}=\frac{64}{3}-\frac{27}{3}=\frac{37}{3}\]